A FATHER IS NOW 7 TIMES AS OLD AS HIS SON. IN 5 YEARS, HE WILL BE 4 TIMES AS OLD AS HIS SON WILL BE THEN. WHAT ARE THEIR PRESENT AGES?
Answer by mathworm7271(5)

Let's call the fathers age f and the sons age s.
Write out the formulas.
F=7S
F+5=4(S+5)
Use the Distributive property on the 2nd equation
F+5=4S+20
Subtract 5 from both sides
F=4S+15
Now substitute
7S=4S+15
Subtract 4S from both sides
3S=15
Divide both sides by three
S=5
Plug in S to the first equation
F=7(5)=35
Therefore, the Father's age is 35 and the Son's age is 7